OpenStudy (anonymous):

The probability that a couple will have two children, so that their first child is a boy and their second child is a girl, is 1/2. True False

OpenStudy (cwrw238):

these are 2 independent events P(boy) first) = 1/2 P(girl second) 1/2 P(boy then girl) = 1/2 * 1/2 = 1/4 independent probabilities are multiplied so false